• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Franklin

    Log entry time 07:44:24 on May20,2010

    Entry number 323469

    keyword=Shift Summary

    Owl Shift Summary
    NO RUNNING
    Shift starts with target pressure going up. Valve closed.

    00:00- vacuum about 1.2E-4
    01:30 vacuum about 1.6E-4 but had a few spikes up to 5 to 8 E-4.
    02:00 spike up to 2.5E-3
    04:30 vacuum about 2.0E-4, spikes up to 5E-4
    06:00 continued spikes, but baseline is still about 2.0E-4